Write down Cp/Cv ratio for monostomic gas


1
Expert's answer
2021-01-24T14:23:54-0500

Answer

Degree of freedom for monoatomic gas for=3.

The ratio of specific heat is given by

"\frac{C_p}{C_v}=\gamma=1+\frac{2}{f}=1+\frac{2}{3}=1.67"